# HoSNNs: Adversarially-Robust Homeostatic Spiking Neural Networks with Adaptive Firing Thresholds

### v3

@@ 3.3.1 Membrane Potential Error Minimization with NDS
Instead of examining the deviation of membrane potential $u_{i}(t|x^{\prime})$ caused by the adversarial input $x^{\prime}$ from the unknown $u_{i}(t|x)$, we define a new error signal $e_{i}(t|x^{\prime})\triangleq u_{i}(t|x^{\prime})-u_{i}^{*}(t|\mathcal{D})$, and replace the optimization problem of (3) by a more practical membrane potential error minimization problem while optimizing the dynamically changing firing threshold $V^{i}_{th}(t|x^{\prime})$ of each neuron $i$:
$\min_{\mathbf{V^{i}_{th}}\in\mathbb{R}^{T}}E_{i}(x^{\prime})=\sum_{t=0}^{T}\frac{1}{2}e_{i}(t|x^{\prime})^{2}\triangleq\sum_{t=0}^{T}\frac{1}{2}(u_{i}(t|x^{\prime})-u_{i}^{*}(t|\mathcal{D}))^{2},$ (6)
@@ 3.3.2 Error Minimization of TA-LIF Neurons as a Second-Order Homeostatic Control System
Subtracting (5) from (1) gives the following dynamics of the error signal $e_{i}(t)$:
$\tau_{m}\frac{de_{i}(t|x^{\prime})}{dt}=-e_{i}(t|x^{\prime})+\Delta I_{i}(t|x^{\prime})-\tau_{m}[r_{i}(x^{\prime})V^{i}_{th}(t|x^{\prime})-r^{*}_{i}(\mathcal{D})V_{th}],$ (7)
@@
where $\varepsilon(t)\triangleq\frac{d\Delta I_{i}(t|x^{\prime})}{dt}$, and we omit the notational dependencies on $x^{\prime}$ for clarity. Importantly, (8) characterizes the error dynamics $e_{i}(t)$ as a second-order control system influenced by the external disturbance $\varepsilon(t)$ with $\frac{dV_{th}^{i}(t)}{dt}$ serving as the control term.
We seek to solve the membrane potential error minimization problem in (6) by designing a control scheme that leads to proper error dynamics based on the second-order error system of (8). To this end, we utilize control signal $\frac{dV_{th}^{i}(t)}{dt}$ to provide a negative feedback control to suppress $e_{i}(t|x^{\prime})$:
@@
Figure 3: Numerical simulations of (8) show that TA-LIF can well suppress the growth of error with time. ($\tau_{m}=r=1$)
TA-LIF model. (1), (2), and (9) together delineate the proposed TA-LIF model. We construct a homeostatic SNN (HoSNN) using TA-LIF neurons, where each TA-LIF neuron maintains its unique $\theta_{i}$ and $V_{th}^{i}(t)$. To extract precise semantic information from $\mathcal{D}$, we collect the NDS for the HoSNN from a well-trained LIF-SNN with identical network configurations. $\theta_{i}$ and network weights $W$ of the HoSNN can be jointly optimized using a training algorithm such as backpropagation.
During inference with the optimized $\theta_{i}$, $V^{i}_{th}(t)$ is adapted in an unsupervised manner according to (9). Intuitively, if TA-LIF neuron $i$ shows abnormal increased activation relative to the reference NDS, $V^{i}_{th}(t)$ would be stepped up to suppress the increase in membrane potential. Conversely, if the neuron is abnormally inhibited, $V^{i}_{th}(t)$ would be tuned down. Figure 3 compares the LIF and TA-LIF models via numerical simulation of (8), showing the growth of error $e_{i}(t)$ over time. We set the $\varepsilon(t)$ as Gaussian white noise $\xi(t)\sim\mathcal{N}(0,1)$ and repeated the simulation 1000 times. The membrane potential error of TA-LIF (red area) is significantly smaller than that of LIF (blue area), revealing TA-LIF’s dynamic robustness under noisy input perturbations.
The feedback control in (9) offers a straightforward means to implement homeostasis, which in turn enhances the adversarial resilience of the proposed HoSNNs. Furthermore, this homeostatic control exhibits two favorable dynamic properties presented next, underscoring its relevance in solving the membrane potential error minimization problem of (6) as a feedback control solution.
@@ 3.3.3 Theoretical Dynamic Properties of TA-LIF Neurons
We highlight key properties of the TA-LIF dynamics from two perspectives: bounded-input bounded-output (BIBO) stability of membrane potential error and suppressed time growth of error in comparison with the standard LIF model. See Appendix A B for a complete derivation of these properties.
@@
If a system is BIBO stable, then the output will be bounded for every input to the system that is bounded. The characteristic equation and its roots of the proposed second-order TA-LIF dynamics (8) incorporating the homeostatic control (9) are:
$\tau_{m}s^{2}+s+r_{i}{\tau_{m}}{\theta_{i}}=0,\quad s_{1,2}=\frac{-1\pm\sqrt{\Delta}}{2\tau_{m}},\quad\Delta=1-4r_{i}{\tau_{m}^{2}}{\theta_{i}}$ (10)
For a second-order system to be BIBO, the roots of its characteristic equation must be a negative real or have a negative real part, which is clearly the case for the TA-LIF model with $\theta_{i}>0$ and $r_{i}>0$, affirming the BIBO stability of the TA-LIF model. This signifies that when the adversarial input perturbation $\varepsilon(t)$ is bounded, the deviation of the TA-LIF neuron’s membrane potential $e_{i}(t)$ is also bounded, demonstrating the good control of error under various attack intensities.
@@
To analyze evolution of membrane potential error induced by injected input perturbations over time, we follow the common practice Gerstner et al. (2014); Abbott and Van Vreeswijk (1993); Brunel (2000); Renart et al. (2004) to approximate $\Delta I(t)$ in (8) as a Wiener process, representing small, independent, and random perturbations. Consequently, the driving force $\varepsilon(t)$ on the right of (8) can be approximated by white noise $\xi(t)$ with zero mean and variance $\sigma^{2}$. By the theory of stochastic differential equations Kloeden et al. (1992), this leads to the following mean square error for the LIF and TA-LIF models, respectively:
$\text{LIF}:\frac{dV^{i}_{th}}{dt}=0\implies\langle e_{i}^{2}(t)\rangle=\frac{\tau_{m}^{2}\sigma^{2}}{2}\left(t-\tau_{m}+\tau_{m}e^{-t/\tau_{m}}\right)=O(\sigma^{2}t)$ (11)
@@ 3.4 Homeostatic SNNs (HoSNNs)
We introduce the homeostatic SNNs (HoSNNs), which deploy TA-LIF neurons as the basic compute units to leverage their noise immunity. Architecturally, HoSNNs can be constructed by adopting typical connectivity such as dense or convolutional layers, with two learnable parameters: synaptic weights $\bm{W}$ and threshold dynamics parameter $\bm{\theta}$ per (9). We extract the network-level NDS $\mathcal{U}_{\text{NET}}(\mathcal{D})$ from a LIF-based SNN with identical architecture well-trained on the clean data distribution $\mathcal{D}$. The HoSNNs optimization problem can be described as:
@@
where $\mathit{x}$ and $\mathit{y}$ are an input/label pair; $\mathcal{L}_{\text{train}}(\cdot)$ is the loss over a training dataset that can include clean/adversarial examples, or a combination of the two; $\mathcal{L}_{\text{mem\_error}}\triangleq\sum_{i=0}^{N}\sum_{t=0}^{T}\frac{1}{2}e_{i}(t|x)^{2}$ is the sum of all neurons’ membrane potential error in (6). In practice, $\mathcal{L}_{\text{mem\_error}}$ is optimized online by the homeostatic control of firing threshold during the forward process. $\mathcal{L}_{\text{train}}(\cdot)$ is optimized by gradient during the backward process, for which any backpropagation based training algorithm such as BPTT Neftci et al. (2019); Wu et al. (2018), BPTR Lee et al. (2020), or TSSL-BP Zhang and Li (2020) can be applied to optimize the network based on (13).
@@ 4.1 Experimental Setup
The proposed HoSNNs are compared with LIF-based SNNs with identical architecture across four benchmark datasets: Fashion-MNIST (FMNIST) Xiao et al. (2017), Street View House Numbers (SVHN) Netzer et al. (2011) CIFAR10 and CIFAR100 Krizhevsky (2009). VGG-like Simonyan and Zisserman (2015) convolutional neural network (CNN) architectures of different sizes and depths are utilized. Widely used methods including FGSM Goodfellow et al. (2014), RFGSM Tramèr et al. (2017), PGD Madry et al. (2017), and BIM Kurakin et al. (2018) are used to generate both white-box and black-box attacks. The dynamically changing firing thresholds of the HoSNNs are exposed to the attacker and utilized in the gradient calculation when generating white-box attacks. We independently trained SNNs with the same architecture and used their white-box attacks as the black-box attacks to the HoSNNs.
@@
For each HoSNN, an LIF-based SNN with an identical architecture is trained on the corresponding clean dataset to derive the NDS. Model training employs the BPTT learning algorithm ($T=5$), leveraging a sigmoid surrogate gradient Xu et al. (2022); Neftci et al. (2019); Wu et al. (2018). The learning rate for each $\theta_{i}$ in (9), which controls the adaptation of firing threshold of TA-LIF neurons, is set to 1/10 of that for the network weights, ensuring hyperparameter stability during training. We ensure that $\theta_{i}$ is non-negative during optimization. We train four types of models: SNNs and HoSNNs on a clean dataset and a weak FGSM-based adversarial training dataset, respectively. For FGSM adversarial training, we set the attack budget to $\epsilon=2/255$ on FMNIST, SVHN and CIFAR10 as in Ding et al. (2022) and $\epsilon=4/255$ on CIFAR100 as in Kundu et al. (2021). For iterative attacks (PGD & BIM), we adopt parameters $\alpha=\epsilon/3$ and $steps=7$ in accordance with Ding et al. (2022). The experimental settings are detailed in the Appendix C.

### v4

@@ 3.3.1 Membrane Potential Error Minimization with NDS
Instead of examining the deviation of membrane potential $u_{i}(t|x^{\prime})$ caused by the adversarial input $x^{\prime}$ from the unknown $u_{i}(t|x)$, we define a new error signal $e_{i}(t|x^{\prime})\triangleq u_{i}(t|x^{\prime})-u_{i}^{*}(t|\mathcal{D})$, and replace the optimization problem of equation 4 by a more practical membrane potential error minimization problem while optimizing the dynamically changing firing threshold $V^{i}_{th}(t|x^{\prime})$ of each neuron $i$:
$\min_{\mathbf{V^{i}_{th}}\in\mathbb{R}^{T}}E_{i}(x^{\prime})=\sum_{t=0}^{T}e_{i}(t|x^{\prime})^{2}\triangleq\sum_{t=0}^{T}(u_{i}(t|x^{\prime})-u_{i}^{*}(t|\mathcal{D}))^{2}$ (7)
@@ 3.3.2 Error Minimization of TA-LIF Neurons as a Second-Order Homeostatic Control System
Subtracting equation 6 from equation 1 gives the following dynamics of the error $e_{i}(t)$:
$\tau_{m}\frac{de_{i}(t|x^{\prime})}{dt}=-e_{i}(t|x^{\prime})+\Delta I_{i}(t|x^{\prime})-\tau_{m}[r_{i}(x^{\prime})V^{i}_{th}(t|x^{\prime})-r^{*}_{i}(\mathcal{D})V_{th}]$ (8)
@@
where $\varepsilon(t)\triangleq\frac{d\Delta I_{i}(t|x^{\prime})}{dt}$, and we omit the notational dependencies on $x^{\prime}$ for clarity. Importantly, equation 9 characterizes the error dynamics $e_{i}(t)$ as a second-order control system influenced by the external disturbance $\varepsilon(t)$ with $\frac{dV_{th}^{i}(t)}{dt}$ serving as the control term.
We seek to solve the membrane potential error minimization problem in equation 7 by designing a control scheme that leads to proper error dynamics based on the second-order error system of equation 9. To this end, we utilize control signal $\frac{dV_{th}^{i}(t)}{dt}$ to provide a negative feedback control to suppress $e_{i}(t|x^{\prime})$:
@@
TA-LIF model.
equation 1, equation 2, and equation 10 together delineate the proposed TA-LIF model. We construct a homeostatic SNN (HoSNN) using TA-LIF neurons, where each TA-LIF neuron maintains its unique $\theta_{i}$ and $V_{th}^{i}(t)$. To extract precise semantic information from $\mathcal{D}$, we collect the NDS for the HoSNN from a well-trained LIF-SNN with identical network configurations. $\theta_{i}$ and network weights $W$ of the HoSNN can be jointly optimized using a training algorithm such as backpropagation.
During inference with the optimized $\theta_{i}$, $V^{i}_{th}(t)$ is adapted in an unsupervised manner according to equation 10. Intuitively, if a TA-LIF neuron $i$ shows a abnormal increased activation relative to the reference NDS, $V^{i}_{th}(t)$ would be stepped up to suppress the increase in membrane potential. Conversely, if the neuron is abnormally inhibited, $V^{i}_{th}(t)$ would be tuned down. Figure 3(a) compares the LIF and TA-LIF models via numerical simulation of equation 9, showing the growth of error $e_{i}(t)$ over time. We set the $\varepsilon(t)$ as Gaussian white noise $\xi(t)\sim\mathcal{N}(0,1)$ and repeated the simulation 1000 times. The membrane potential error of TA-LIF (red area) is significantly smaller than that of LIF (blue area), revealing TA-LIF’s dynamic robustness under noisy input perturbations.
The feedback control in equation 10 offers a straightforward means to implement homeostasis, which in turn enhances the adversarial resilience of the proposed HoSNNs. Furthermore, this homeostatic control exhibits two favorable dynamic properties presented next, underscoring its relevance in solving the membrane potential error minimization problem of equation 7 as a feedback control solution.
(a) Numerical simulations of equation 9, with parameters $\tau_{m}=1,\;r=1$.
(b) Box plot of the post-synaptic current relative error distribution per layer of a SNN and a HoSNN.
Figure 3: (a) Numerical simulations of equation 9 show that TA-LIF can well suppress the growth of error with time. (b) Box plot of the post-synaptic current relative error distribution per layer of a SNN and a HoSNN, trained with FGSM adversarial training with $\epsilon=2/255$ and attacked by a same CIFAR10 black-box PGD7 dataset with $\epsilon=8/255$.
@@ 3.3.3 Theoretical Dynamic Properties of TA-LIF Neurons
We highlight key properties of the TA-LIF dynamics from two perspectives: bounded-input bounded-output (BIBO) stability of membrane potential error and suppressed time growth of error in comparison with the standard LIF model. See Appendix A for complete derivation of these properties.
@@
If a system is BIBO stable, then the output will be bounded for every input to the system that is bounded. The characteristic equation and its roots of the proposed second-order TA-LIF dynamics equation 9 incorporating the homeostatic control equation 10 are:
$\tau_{m}s^{2}+s+r_{i}{\tau_{m}}{\theta_{i}}=0,\ s_{1,2}=\frac{-1\pm\sqrt{1-4r_{i}{\tau_{m}^{2}}{\theta_{i}}}}{2\tau_{m}}$ (11)
For a second-order system to be BIBO, the roots of its characteristic equation must be a negative real or have a negative real part, which is clearly the case for the TA-LIF model with both $\theta_{i}$ and $r_{i}>0$, affirming the BIBO stability of the TA-LIF model. This signifies that when the adversarial input perturbation $\varepsilon(t)$ is bounded, the deviation of the TA-LIF neuron’s membrane potential $e_{i}(t)$ is also bounded, showing the good control of error under various attack intensities.
@@
To analyze the evolution of membrane potential error induced by injected input perturbations over time, we follow the common practice (Gerstner et al., 2014; Abbott & Van Vreeswijk, 1993; Brunel, 2000; Renart et al., 2004) to approximate $\Delta I(t)$ in equation 9 as a Wiener process, representing small, independent, and random perturbations. Consequently, the driving force $\varepsilon(t)$ on the right of equation 9 can be approximated by white noise $\xi(t)$ with zero mean and variance $\sigma^{2}$. By the theory of stochastic differential equations (Kloeden et al., 1992), this leads to the following mean square error for the LIF and TA-LIF models, respectively:
$\text{LIF}:\frac{dV^{i}_{th}}{dt}=0\implies\langle e_{i}^{2}(t)\rangle=O(\sigma^{2}t)=\frac{\tau_{m}^{2}\sigma^{2}}{2}\left(t-\tau_{m}+\tau_{m}e^{-t/\tau_{m}}\right)$ (12)
@@ 3.4 Homeostatic SNNs (HoSNNs)
We further introduce the homeostatic SNNs (HoSNNs), which deploy TA-LIF neurons as the basic compute units to leverage their noise immunity. Architecturally, HoSNNs can be constructed by adopting typical connectivity such as dense or convolutional layers, with two learnable parameters: synaptic weights $\bm{W}$ and threshold dynamics parameter $\bm{\theta}$ per equation 10. We extract the network-level NDS $\mathcal{U}_{\text{NET}}(\mathcal{D})$ from a LIF-based SNN with identical architecture well-trained on the clean data distribution $\mathcal{D}$. The HoSNNs optimization problem can be described as:
@@
where $\mathit{x}$ and $\mathit{y}$ are an input/label pair; $\mathcal{L}_{\text{train}}(\cdot)$ is the loss over a training dataset that can include clean/adversarial examples, or a combination of the two; $\mathcal{L}_{\text{mem}}\triangleq\sum_{i=0}^{N}\sum_{t=0}^{T}e_{i}(t|x)^{2}$ is the sum of all neurons’ membrane potential error in equation 7. In practice, $\mathcal{L}_{\text{mem}}$ is optimized online by the homeostatic control of firing threshold during the forward process. $\mathcal{L}_{\text{train}}(\cdot)$ is optimized by gradient during the backward process, for which any backpropagation based training algorithm such as BPTT (Neftci et al., 2019; Wu et al., 2018), BPTR (Lee et al., 2020), or TSSL-BP (Zhang & Li, 2020) can be applied to optimize the network based on equation 14.
3.5 Complexity Analysis of HoSNN
In a standard Spiking Neural Network (SNN), the number of neurons is $n$, with weights $w=O(n^{2})$ and training time $T$. For Higher-order Spiking Neural Networks (HoSNNs), additional space is required to store the learnable parameters $\theta_{i}=O(n)$, but this storage cost is negligible compared to the network weights. The primary additional cost in training a HoSNN is training the baseline SNN first. The extra time due to the TA-LIF model’s dynamic threshold calculation is minimal (<10%). Therefore, the overall training time for a HoSNN is about $2T$. The dynamic threshold adjustment during inference has negligible impact, so the inference cost of a trained HoSNN is almost identical to that of an SNN. There are some challenges in the training of HoSNN. (1) $\theta_{i}$ Initialization: Initializing $\theta_{i}$ too large (> 1) can destabilize training. It is recommended to set $\theta_{i}$ between 0 and 0.5, with smaller values needed for complex datasets. (2) Optimizer Choice: HoSNNs converge slower in later training stages, especially in adversarial settings. Using the Adam optimizer with a cosine decay learning rate helps achieve similar convergence speeds as SNNs.
@@ 4.1 Experimental Setup
The proposed HoSNNs are compared with LIF-based SNNs with identical architecture across four benchmark datasets: Fashion-MNIST (FMNIST) (Xiao et al., 2017), Street View House Numbers (SVHN) (Netzer et al., 2011) CIFAR10 and CIFAR100 (Krizhevsky, 2009). VGG-5/9/11 (Simonyan & Zisserman, 2015) convolutional neural network (CNN) architectures of different sizes and depths are utilized. Widely used methods including FGSM (Goodfellow et al., 2014), RFGSM (Tramèr et al., 2017), PGD (iteration = 7, 20, 40) (Madry et al., 2017), and BIM (Kurakin et al., 2018) are used to generate both white-box and black-box attacks. To test stronger attacks, we introduce APGD (Croce & Hein, 2020a) in our white-box attack, with cross entropy loss ($\text{APGD}_{\text{CE}}$), difference of logits ratio loss ($\text{APGD}_{\text{DLR}}$) and targeted attack (T-APGD). Unless otherwise specified, PGD refers to PGD7. The dynamically changing firing thresholds of the HoSNNs are exposed to the attacker and utilized in the gradient calculation when generating white-box attacks. We independently trained SNNs with the same architecture and used their white-box attacks as the black-box attacks to the HoSNNs.
@@
For each HoSNN, an LIF-based SNN with an identical architecture is trained on the corresponding clean dataset to derive the NDS. Model training employs the BPTT learning algorithm ($T=5$), leveraging a sigmoid surrogate gradient (Xu et al., 2022; Neftci et al., 2019; Wu et al., 2018). The learning rate for each $\theta_{i}$ in equation 10, which controls the adaptation of the firing threshold of TA-LIF neurons, is set to 1/10 of that for the network weights, ensuring hyperparameter stability during training. We ensure that $\theta_{i}$ is non-negative during optimization. We train four types of models: SNNs and HoSNNs on a clean dataset and a weak FGSM-based adversarial training dataset, respectively. For FGSM adversarial training, we set the attack budget to $\epsilon=2/255$ on FMNIST, SVHN and CIFAR10 as in (Ding et al., 2022) and $\epsilon=4/255$ on CIFAR100 as in (Kundu et al., 2021). For iterative attacks (PGD & BIM), we adopt parameters $\alpha=2.5*\epsilon/steps$ and $steps=7,20,40$ in accordance with (Ding et al., 2022). Mode experimental settings are in the Appendix C.
4.2 Adversarial Robustness under White-box Attacks
Robustness without adversarial training
Table 1 compares HoSNNs and SNNs intrinsic resilience without adversarial training. Both types of network are trained exclusively on the clean dataset, and then subjected to $\epsilon=8/255$ white-box adversarial attacks. The HoSNNs show consistently higher accuracy than the SNN counterparts under all four datasets and four attacks. For example, on CIFAR-10, the HoSNN significantly improves accuracy from 20.86% to 54.76% under FGSM, from 0.54% to 15.32% under PGD7, from 0.69% to 10.35% under APGD with cross entropy loss, from 4.44% to 16.02% under T-APGD.
Robustness with adversarial training
In Table 2, we evaluate the enhanced robustness of HoSNNs under adversarial training. We train both the SNNs and HoSNNs using FGSM adversarial training and then expose them to $\epsilon=8/255$ white-box attacks. The results show a significant boost in robustness for HoSNNs when using adversarial training. Furthermore, the HoSNNs noticeably outperform the SNNs trained using the same FGSM adversarial training. For example, on CIFAR10, the HoSNNs improve the accuracy of the corresponding SNN to 63.98% from 37.93% under FGSM attack, and to 43.33% from 12.42% under PGD7 attack, from 8.23% to 38.89% under APGD-CE attack. On CIFAR100, the HoSNN improves the accuracy to 16.66% from 8.82% under PGD7 attack, from 7.75% to 12.55% under APGD-CE attack.
Dataset Net Clean FGSM RFGSM BIM7 PGD7 PGD20 PGD40 $\text{APGD}_{\text{CE}}$ $\text{APGD}_{\text{DLR}}$ T-APGD
Fashion MNIST $\times$ 92.92 56.01 70.02 38.85 30.90 28.73 27.88 23.22 40.25 39.67
$\checkmark$ $\mathbf{92.96}$ $\mathbf{65.36}$ $\mathbf{76.10}$ $\mathbf{49.02}$ $\mathbf{35.79}$ $\mathbf{32.99}$ $\mathbf{32.63}$ $\mathbf{37.78}$ $\mathbf{61.01}$ $\mathbf{50.57}$
SVHN $\times$ $\mathbf{95.51}$ 26.07 42.94 2.26 0.44 0.18 0.13 0.11 7.73 0.81
$\checkmark$ 93.55 $\mathbf{44.87}$ $\mathbf{57.27}$ $\mathbf{12.91}$ $\mathbf{4.33}$ $\mathbf{1.87}$ $\mathbf{1.47}$ $\mathbf{3.09}$ $\mathbf{12.26}$ $\mathbf{5.32}$
CIFAR 10 $\times$ $\mathbf{92.47}$ 20.86 38.72 3.29 0.54 0.38 0.3 0.69 7.03 4.44
$\checkmark$ 92.43 $\mathbf{54.76}$ $\mathbf{62.33}$ $\mathbf{28.06}$ $\mathbf{15.32}$ $\mathbf{11.58}$ $\mathbf{10.74}$ $\mathbf{10.35}$ $\mathbf{27.39}$ $\mathbf{16.02}$
CIFAR 100 $\times$ $\mathbf{74.00}$ 5.74 8.94 0.10 0.04 0.01 0.00 0.00 0.01 0.01
$\checkmark$ 71.98 $\mathbf{13.48}$ $\mathbf{12.27}$ $\mathbf{0.50}$ $\mathbf{0.19}$ $\mathbf{0.02}$ $\mathbf{0.02}$ $\mathbf{2.55}$ $\mathbf{0.02}$ $\mathbf{0.02}$
Table 1: Training on clean dataset. Whitebox attack results on Fashion-MNIST, SVHN, CIFAR10 and CIFAR100 under various types of attack with an intensity of $\epsilon=8/255$. The data on the left and right are based on training using the clean and weak FGSM datasets, respectively. HoSNNs (denoted as $\checkmark$) provide greater robustness than SNNs (denoted as $\times$) under all attacks and datasets.
Dataset Net Clean FGSM RFGSM BIM7 PGD7 PGD20 PGD40 $\text{APGD}_{\text{CE}}$ $\text{APGD}_{\text{DLR}}$ T-APGD
Fashion MNIST $\times$ 92.08 74.12 83.12 68.36 62.92 61.97 61.43 55.47 68.25 62.56
$\checkmark$ $\mathbf{92.31}$ $\mathbf{84.7}$ $\mathbf{87.99}$ $\mathbf{79.61}$ $\mathbf{74.91}$ $\mathbf{73.53}$ $\mathbf{73.21}$ $\mathbf{67.94}$ $\mathbf{76.24}$ $\mathbf{74.48}$
SVHN $\times$ $\mathbf{93.85}$ 48.37 68.81 30.09 18.46 15.52 14.79 9.70 25.66 20.82
$\checkmark$ 92.84 $\mathbf{61.78}$ $\mathbf{75.60}$ $\mathbf{48.83}$ $\mathbf{36.82}$ $\mathbf{32.24}$ $\mathbf{30.89}$ $\mathbf{28.20}$ $\mathbf{47.89}$ $\mathbf{43.96}$
CIFAR 10 $\times$ $\mathbf{91.87}$ 37.93 59.50 22.31 12.42 11.03 10.63 8.23 16.8 14.62
$\checkmark$ 90.00 $\mathbf{63.98}$ $\mathbf{71.07}$ $\mathbf{52.33}$ $\mathbf{43.33}$ $\mathbf{40.97}$ $\mathbf{40.02}$ $\mathbf{38.89}$ $\mathbf{37.94}$ $\mathbf{41.69}$
CIFAR 100 $\times$ $\mathbf{68.72}$ 22.54 36.93 13.58 8.82 7.88 7.52 7.75 7.51 5.4
$\checkmark$ 64.64 $\mathbf{26.97}$ $\mathbf{41.45}$ $\mathbf{21.09}$ $\mathbf{16.66}$ $\mathbf{15.83}$ $\mathbf{15.37}$ $\mathbf{12.55}$ $\mathbf{13.66}$ $\mathbf{10.2}$
Table 2: Training on FGSM dataset ($\epsilon=2/255$). Whitebox attack results on Fashion-MNIST, SVHN, CIFAR10 and CIFAR100 under various types of attack with an intensity of $\epsilon=8/255$. The data on the left and right are based on training using the clean and weak FGSM datasets, respectively. HoSNNs (denoted as $\checkmark$) provide greater robustness than SNNs (denoted as $\times$) under all attacks and datasets.
4.3 Adversarial Robustness under Black-box Attacks
Layer-wise Relative Error of Post-synaptic Currents
Perturbation in membrane potential caused by adversarial inputs can alter the output spike train of each neuron, and the resulting shifts in its post-synaptic current (PSC) can propagate through the successive layers. The PSC $a_{i}(t)$ is calculated by equation 3. To reveal the source of HoSNNs robustness, we examine the relative error in PSC induced by adversarial attacks. On CIFAR10, we use the $\epsilon=8/255$ black-box PGD7 attack to attack SNN and HoSNN trained with the $\epsilon=2/255$ FGSM adversarial training. For neuron $i$, we record its PSC $a_{i}(t)$ and $a^{\prime}_{i}(t)$ under each pair of clean and adversarial input, respectively, and then calculate the difference of $e^{PSC}_{i}(t)\triangleq|a_{i}(t)-a^{\prime}_{i}(t)|$ as an error metric. The layer-wise distributions of the relative error $e^{PSC}_{i}$ are plotted in Figure 3(b) and Figure 4.
The experiment results show that each error distribution of the HoSNN has a significantly reduced mean compared with that of the SNN, and has its probability mass concentrated on low PSC error values, revealing the favorable internal self-stabilization introduced by the proposed homeostasis. Correspondingly, the HoSNN delivers an accuracy of 76.62%, significantly surpassing the 46.97% accuracy of the SNN.
Figure 4: Distribution of post-synaptic current relative error of a SNN and HoSNN trained with FGSM adversarial training with $\epsilon=2/255$ and attacked by black-box PGD7 attack with $\epsilon=8/255$.
Black Box Robustness
Dataset Net Clean FGSM RFGSM PGD7 BIM7
Fashion MNIST $\times$ 92.08 66.26 80.09 74.08 73.43
$\checkmark$ $\mathbf{92.31}$ $\mathbf{68.31}$ $\mathbf{80.73}$ $\mathbf{75.12}$ $\mathbf{74.09}$
SVHN $\times$ $\mathbf{93.85}$ 17.37 42.64 21.16 36.70
$\checkmark$ 92.84 $\mathbf{19.08}$ $\mathbf{44.57}$ $\mathbf{25.97}$ $\mathbf{40.75}$
CIFAR 10 $\times$ $\mathbf{91.87}$ 13.48 8.79 0.11 0.31
$\checkmark$ 90.00 $\mathbf{25.18}$ $\mathbf{31.11}$ $\mathbf{13.42}$ $\mathbf{23.34}$
CIFAR 100 $\times$ $\mathbf{68.72}$ 12.18 17.87 6.74 17.36
$\checkmark$ 64.64 $\mathbf{14.54}$ $\mathbf{24.32}$ $\mathbf{16.90}$ $\mathbf{32.04}$
Table 3: SNNs and HoSNNs black-box attack accuracy, trained with FGSM adversarial training and tested by different black-box attack methods with $\epsilon=32/255$.
Table 3 evaluates the robustness of SNNs (denoted as $\times$) and HoSNNs (denoted as $\checkmark$) against black-box attacks. All models are trained by weak FGSM adversarial training and tested by $\epsilon=32/255$ black-box attacks generated using separately trained SNNs with identical architecture. Table 3 shows that the HoSNNs exhibit significantly stronger black-box robustness than the SNN counterparts. For example on CIFAR-10, the HoSNN outperforms the traditional SNN under the FGSM and PGD7 attacks with 11.7% and 13.31% accuracy improvements, respectively. On CIFAR-100, HoSNN improves the PGD7 accuracy from 6.74% to 16.90%.
4.4 Checklist for gradient obfuscation
For a detailed check of gradient obfuscation (Athalye et al., 2018; Carlini et al., 2019), we provide comprehensive inspection and data (in Appendix D). HoSNNs pass all five tests as show in Table 4
Items to identify gradient obfuscation HoSNN Experiment
(1) Single-step attack performs better compared to iterative attacks Fig 5 and Table 8
(2) Black-box attacks perform better compared to white-box attacks Fig 6 and Table 9
(3) Increasing perturbation bound can’t increase attack strength Fig 7 and Table 10
(4) Unbounded attacks can’t reach $100\%$ success Fig 7 and Table 10
(5) Adversarial example can be found through random sampling Fig 8
Table 4: Checklist for gradient obfuscation
For Test (1)
We plot the curves of white-box FGSM and PGD7 attacks on four datasets in Figure 5 and Table 8, with attack budgets $\epsilon$ from 0 to 64/255 to ensure that the network can be completely fooled. The red curve is the accuracy under FGSM, while the blue curve is PGD7. From the Figure 5 we can confirm that all iterative attacks are much stronger than single-step attack.